# High-Quality Spatial Reconstruction and Orthoimage Generation Using Efficient 2D Gaussian Splatting

### v1

@@ Abstract
I INTRODUCTION
@@
Digital Orthophoto Maps (DOMs) from Unmanned Aerial Vehicle (UAV) surveying provide dense textural and geometric data, useful in fields like agriculture [1], environmental monitoring [2], and disaster assessment [3]. DOMs are typically created by combining a Digital Elevation Model (DEM) with images captured from a fixed perspective, resulting in a nadir image of the target surface to correct projection distortion due to terrain and oblique photography [4]. Yet, facade occlusions of object can produce artifacts and incorrect geometry. True Digital Orthophoto Maps (TDOMs) address this by incorporating Digital Surface Models (DSM) and using visibility checks to reduce occlusion effects and enhance map accuracy [5].
@@
Over the past decades, numerous methods have been suggested for the generation of True Digital Orthophoto Maps (TDOMs). The Z-buffer method [6, 7, 8], for example, resolves visibility by keeping a note of distances between object points and perspective center corresponding to image pixels and selecting the nearest point. Another method identifies occlusions from the angle between the perspective center and Digital Surface Model (DSM) units with adaptive radial and spiral sweeps, and according to the height analysis of the light rays [9]. In recent years, few learning-based methods have been used to generate TDOMs, but have the drawbacks of limited generalization and high dependence on LiDAR data. For instance, Ebrahimikia and Hosseininaveh’s method[10] identifies building edges in images and alters the DSM, whereas other methods directly produce TDOMs through the use of Generative Adversarial Networks (GANs)[11, 12].
@@
Neural Radiance Fields (NeRF)[13, 14] and 3D Gaussian Splatting (3DGS) [15], new rendering-based methods, have introduced new solutions to the development of True Digital Orthophoto Maps (TDOMs). NeRF is made more efficient with implicit 3D scene representation and differentiable rendering, further aided by approaches such as Plenoxels [16] and Ortho-NeRF [17]. However, NeRF is still troubled by long training times and heavy rendering loads that hinder its real-time implementation. Conversely, 3DGS uses Gaussian kernels to encode 3D scene data explicitly and accelerates training and rendering through parallel processing, e.g., TOrtho-Gaussian [18], at more than 100 frames per second. The scalability limitations of 3DGS in large-scale scene primarily stem from the hardware-dependent constraint that the number of 3D Gaussian kernels is bounded by the video memory capacity of graphics processing units. Furthermore, fidelity needs to be prioritized, particularly when dealing with reflections and slender structures, where blurring and aliasing may occur.
@@
This work describes a new TDOM-generation method based on a variant of 3DGS called 2DGS [19]. Surface reconstruction quality and perspective consistency are essential for TDOM generation. 2DGS resolves the perspective inconsistency of 3DGS by enhancing both geometric representation and rendering mechanisms. By utilizing normal information to refine surface structures, 2DGS yields higher-quality 3D reconstructions. Many TDOM applications require depth information to accurately represent the 3D structure of the scene, and our method addresses this by extracting the corresponding depth maps during the rendering process, thereby enhancing the overall reconstruction accuracy. Inspired by VastGaussian [20] and CityGaussian [21], we adopt a divide-and-conquer strategy to enable large-scale 2DGS training and high-resolution TDOM rendering with limited computational resources, as demonstrated in LABEL:fig:presentation.
In summary, we make the following contributions:
  1. 1.
    We propose a TDOM-generation framework based on a highly efficient differentiable 2D Gaussian renderer, which ensures geometric fidelity and consistency across multi-view observations.
  2. 2.
    We introduce a geometry-aware depth estimation module that directly derives depth maps during the rendering process.
  3. 3.
    Our approach leverages a divide-and-conquer strategy to enable efficient large-scene reconstruction and high-resolution TDOM rendering under constrained computational resources, bridging the gap between quality and practicality.
II RELATED WORKS
II-A Traditional TDOM Generation Methods
@@
The Z-buffer method [6] is one of the earliest proposed visibility analysis algorithms, characterized by its simplicity in principle, but exhibits strong dependence on a high-precision DSM. Subsequently, the angle-based occlusion detection algorithm [9] was proposed, which identified occlusion by sequentially analyzing the angles of projected rays along radial directions in the DSM. Inspired by ray tracing algorithms, Wang et al. [22] optimized the algorithm to ensure smooth and stable edge detection while reducing computation time. Alternative approaches include the height-based shadow detection methodology initially developed by Bang [23]. This method identifies occlusion by comparing the height of projected rays with the DSM height at regular horizontal intervals. There is also a vector polygon-based algorithm [24, 25], which projects the vectors polygons representing building surfaces into the image space and evaluates occupancy priority in overlapping areas to acquire the actual coverage relationship. Wang et al. [26] introduced a method for enhancing shadow information in color aerial imagery and proposed an information lost area compensation approach based on texture matching.
Each of these methods has distinct characteristics and specific application scenarios. However, they also come with inherent limitations. For instance, they rely heavily on high-quality DSM or DBM, the acquisition of which is often costly. Moreover, these methods still face challenges in achieving precise edge detection [27], natural texture compensation [12], and a streamlined, efficient workflow [28, 29] . These difficulties become even more pronounced when generating large-scale TDOM [30].
II-B Deep Learning-based Methods
Recent methodological advancements have increasingly adopted deep learning frameworks to replace some intermediate steps in traditional TDOM generation workflows [12, 31, 32]. For example, Urban-SnowflakeNet [32] utilizes a deep learning network to extract features from preprocessed rooftop point clouds. It reconstructs building point clouds to aid in DSM rectification and TDOM generation, thereby effectively reducing distortions at building edges in urban scenes. However, these methods still rely on a combination of vision and LiDAR, requiring high-quality LiDAR point clouds and potentially failing in certain scenarios.
@@ II-B Deep Learning-based Methods
II-C Differentiable Rendering-based Methods
The new view synthesis method based on differentiable volume rendering has significantly improved the quality of reconstructed images, providing a novel approach for pure vision-based TDOM generation. By integrating differentiable volume rendering, NeRF[33] trains an implicit neural radiance field, achieving remarkable-fidelity and highly continuous scene reconstruction. Many speed-optimized variants of NeRF, such as Instant-NGP [34] and Plenoxels [16], have been applied to TDOM rendering, achieving image quality on par with traditional approaches [35, 17].
@@ II-C Differentiable Rendering-based Methods
III METHOD
The entire TDOM production process is shown in Fig 1. First, we initiate the point cloud as 2D Gaussians with original 2DGS training strategy presented in Section III-A. The orthographic camera model and the depth map generation method are described respectively in Section III-B and Section III-C. For the sake of training large-scale scenes, we employ a divide-and-conquer approach, and the seamlessly merged 2D Gaussians are rendered using Batch Rasterization to obtain high-resolution TDOM ,which is detailed in Section III-D.
III-A Preliminaries
@@
2DGS proposes flattening 3D Gaussian ellipsoids into 2D Gaussian elliptical disks. The planar-like shape of 2D Gaussian not only aligns more closely with the needs of surface reconstruction but also resolves the inconsistency in depth perception in 3DGS. As illustrated in Fig 2, this inconsistency arises when we observe Gaussian ellipsoids from different views, and the actual observed plane also varies. 2DGS constructs a local coordinate system for each 2D Gaussian. Any point $\mathbf{P}(u,v)$ in its local coordinate system can be transformed into its corresponding world coordinates with Gaussian center $\mu$, two principal tangential vectors $(\mathbf{t}_{u},\mathbf{t}_{v})$ and the scaling factors $(s_{u},s_{v})$:
@@ III-B Orthographic Projection of 2DGS
To obtain an orthographic image, we modify the projection method for the centers of the Gaussian primitives to orthographic projection. Shown as Fig 4, in the orthographic camera model, all rays are parallel to the optical axis of the camera. These two camera models are both demonstrated in Fig 3. Orthographic projection defines a square frustum as the visible space, and the projection matrix is defined as follows:
@@
III-C Orthographic Depth Maps
@@
Since 2D Gaussian primitives are elliptical disks, when observed from a specific viewpoint, the primitives can degenerate from surfaces into lines, which is a degeneracy issue present in 2DGS. Shown as Fig 4, When the Gaussian value $G_{i}(\mathbf{x})$ at the ray-splat intersection is too small, we create a standard 2D Gaussian distribution $g_{i}^{2D}(\mathbf{x})$, using the projection of the Gaussian primitive center on the image plane as the mean. By comparing the values at the intersections of the ray with the two Gaussians, when $g_{i}^{2D}(\mathbf{x})$ is greater than $G_{i}(\mathbf{x})$, the 2D Gaussian for the current view is considered degenerate, and $d_{i}$ is taken as the z-axis coordinate of the Gaussian center in view space.
III-D High-Resolution TDOM Rendering Strategy
@@ IV EXPERIMENTS
IV-A Experiments setup
Dataset: We selected several scenes with varying scales and characteristics from the NPU DroneMap Dataset[38], where the largest scene covers an area of 1.958 km². The chosen scenes include diverse land use types, including urban high-rise buildings, rural cottages, farmland, highways, and water bodies. The dataset was created by Bu et al., recorded using a custom-built hexacopter equipped with a GoPro Hero3+ camera and a DJI Phantom3. Each image has a 1920×1080 resolution. Moreover, we have also chosen a sample dataset provided by Agisoft [39], recorded by Canon DIGITAL IXUS 120 IS, which encompasses ground control point (GCP) information, to conduct a quantitative assessment of the TDOM’s quality. The panoramic images generated by our method are shown in Fig 5.
Implementation: The initial point cloud and camera poses obtained from COLMAP sparse reconstruction first undergo Manhattan alignment to align the scene’s bounding box with the coordinate axes. Each scene is divided into 8 partitions following the progressive partition strategy. Training is conducted in two batches across 4 NVIDIA GeForce RTX 3090 GPUs for 60,000 iterations, during this period, the attributes of the 2D Gaussians are continuously optimized. After training, the gaussians are filtered according to their central positions using bounding boxes from camera-based scene partitioning. The average training time for a scene is around 7 hours. Soon, the processed points will be stitched and deduplicated for the subsequent rendering. We use the orthographic camera model mentioned in Section III-B to render orthophotos. The complete image is divided into 2$\times$2 tiles and rendered one by one. When each pixel is rendered in the GPU unit, the orthographic depth is also calculated.
TABLE I: Absolute distance error comparison between Metashape, ContextCapture, Pix4Dmapper and our method.
Pairs GT(m) Aligned Distance(m) Absolute Error(m)$\downarrow$
Metashape ContextCapture Pix4Dmapper Ours Metashape ContextCapture Pix4Dmapper Ours
101-202 41.226 40.348 40.625 40.684 40.510 0.876 0.601 0.542 0.716
102-203 73.158 72.933 72.617 72.804 73.875 0.942 1.257 1.071 0.717
202-205 133.702 132.651 131.994 132.288 132.683 1.708 1.414 1.019 1.057
203-206 119.768 120.088 118.796 118.423 119.456 0.968 1.339 0.306 0.326
102-206 85.800 86.067 84.658 84.697 86.173 1.142 1.103 0.373 0.267
202-206 121.098 123.834 122.349 122.554 123.884 1.251 1.456 2.786 2.736
IV-B TDOM Evaluation
@@
  • Buildings: In terms of high-quality TDOM, the buildings’ edges must be continuous and sharp, and their facades must not be visible. Our method excellently reconstructs the building boundaries and their occlusion relationships. In Fig 6(a) and Fig 6(b), it can be observed that other methods generally exist visible building facades, with Map2DFusion showing the most severe issues due to direct photo mosaicking without image correction. Additionally, ContextCapture has blurred edges, Metashape displays noticeable irregular boundaries, and Pix4Dmapper suffers from misaligned building edges.
  • Thin structures: Fig 6(c) and Fig 6(d) respectively demonstrate the restoration effects of the power tower and the air conditioner guardrail. Metashape consistently blurs and granulates scene information; the supports of the power tower become thickened and distorted in Map2DFusion, and are diminished in ContextCapture. The railings of the guardrail are significantly misaligned in all four software programs. In our method, these thin structures are clearer, and the details are well restored even under conditions of strong light and shadow contrast.
In the quantitative evaluation, the absolute distance error between two GCPs is applied as a criterion to evaluate the accuracy of the TDOM. We calculate the distances between the GCPs using both their true coordinate values and the coordinates extracted from the local coordinate system of the TDOM, while the true horizontal distances between the GCPs are computed using the Haversine formula. By aligning one pair of GCPs, we obtain a scaling factor, and then calculate the error between the aligned GCP distances from the TDOM and their true values. We compare our results with those from three commercial software solutions: Metashape, ContextCapture, and Pix4Dmapper. All the methods aligned the pair of control points GCP-101 and GCP-102. In our method, the minimum absolute error for GCP pairs is 0.267m, and the maximum absolute error is 2.736m. Our method maintains errors within 1.0m in the majority of cases and demonstrates comparable or superior accuracy relative to commercial software across most point pairs. Specific data can be found in Table I.
@@ IV-C Depth Map Evaluation
In the qualitative experiments, we replace the Red channel of the orthophoto with the depth map to facilitate an intuitive visualization of the correspondence between the depth information and the image. Additionally, we use the Canny edge detection algorithm [43] to extract building edges from the depth map and overlay them onto our generated TDOM to observe the alignment of the edges. As illustrated in Fig 7(1), The deeper red hue in the pixel indicates the higher height of the relevant surface object relative to the ground. The red areas predominantly cover the buildings on the surface, which aligns with real-world logic. In Fig 7(2), the edges extracted from the depth map distinctly outline the boundaries of the structures within the TDOM, as well as some of the taller trees on the terrain. Our depth map adds a third dimension of height information to the TDOM, which is beneficial for subsequent tasks such as image detection and semantic segmentation in TDOM.
V CONCLUSIONS

### v2

@@ Abstract
keywords:
True Digital Orthophoto Map (TDOM), Gaussian Splatting, Spatial Reconstruction, Occlusion Detection.
1 Introduction
@@
Digital Orthophoto Maps (DOMs) from Unmanned Aerial Vehicle (UAV) surveying provide dense textural and geometric data, useful in fields like agriculture, environmental monitoring, and disaster assessment. DOMs are typically created by combining a Digital Elevation Model (DEM) with images captured from a fixed perspective, resulting in a nadir image of the target surface to correct projection distortion due to terrain and oblique photography [1]. Yet, facade occlusions of object can produce artifacts and incorrect geometry. True Digital Orthophoto Maps (TDOMs) address this by incorporating Digital Surface Models (DSM) and using visibility checks to reduce occlusion effects and enhance map accuracy [2].
@@
Over the past decades, numerous methods have been suggested for the generation of True Digital Orthophoto Maps (TDOMs). The Z-buffer method [3], for example, resolves visibility by keeping a note of distances between object points and perspective center corresponding to image pixels and selecting the nearest point. Another method identifies occlusions from the angle between the perspective center and Digital Surface Model (DSM) units with adaptive radial and spiral sweeps, and according to the height analysis of the light rays [4]. In recent years, few learning-based methods have been used to generate TDOMs, but have the drawbacks of limited generalization and high dependence on LiDAR data. For instance, Ebrahimikia and Hosseininaveh’s method[5] identifies building edges in images and alters the DSM, whereas other methods directly produce TDOMs through the use of Generative Adversarial Networks (GANs)[6].
@@
Neural Radiance Fields (NeRF)[7] and 3D Gaussian Splatting (3DGS) [8], new rendering-based methods, have introduced new solutions to the development of True Digital Orthophoto Maps (TDOMs). NeRF is made more efficient with implicit 3D scene representation and differentiable rendering, further aided by approaches such as Ortho-NeRF [9]. However, NeRF is still troubled by long training times and heavy rendering loads that hinder its real-time implementation. Conversely, 3DGS uses Gaussian kernels to encode 3D scene data explicitly and accelerates training and rendering through parallel processing, e.g., TOrtho-Gaussian [10], at more than 100 frames per second. The scalability limitations of 3DGS in large-scale scene primarily stem from the hardware-dependent constraint that the number of 3D Gaussian kernels is bounded by the video memory capacity of graphics processing units. Furthermore, fidelity needs to be prioritized, particularly when dealing with reflections and slender structures, where blurring and aliasing may occur.
@@
This work describes a new TDOM-generation method based on a variant of 3DGS called 2DGS [11]. Surface reconstruction quality and perspective consistency are essential for TDOM generation. 2DGS resolves the perspective inconsistency of 3DGS by enhancing both geometric representation and rendering mechanisms. By utilizing normal information to refine surface structures, 2DGS yields higher-quality 3D reconstructions. Many TDOM applications require depth information to accurately represent the 3D structure of the scene, and our method addresses this by extracting the corresponding depth maps during the rendering process, thereby enhancing the overall reconstruction accuracy. Inspired by VastGaussian [12], we adopt a divide-and-conquer strategy to enable large-scale 2DGS training and high-resolution TDOM rendering with limited computational resources.
2 Related Works
2.1 Traditional TDOM Generation Methods
@@
The Z-buffer method [3] is one of the earliest proposed visibility analysis algorithms, characterized by its simplicity in principle, but exhibits strong dependence on a high-precision DSM. Subsequently, the angle-based occlusion detection algorithm [4] was proposed, which identified occlusion by sequentially analyzing the angles of projected rays along radial directions in the DSM. Inspired by ray tracing algorithms, Wang et al. [13] optimized the algorithm to ensure smooth and stable edge detection while reducing computation time. There is also a vector polygon-based algorithm [14], which projects the vectors polygons representing building surfaces into the image space and evaluates occupancy priority in overlapping areas to acquire the actual coverage relationship.
Each of these methods has distinct characteristics and specific application scenarios. However, they also come with inherent limitations. For instance, they rely heavily on high-quality DSM or DBM, the acquisition of which is often costly. Moreover, these methods still face challenges in achieving precise edge detection, natural texture compensation [6], and a streamlined, efficient workflow [15] . These difficulties become even more pronounced when generating large-scale TDOM.
2.2 Deep Learning-based Methods
Recent methodological advancements have increasingly adopted deep learning frameworks to replace some intermediate steps in traditional TDOM generation workflows [6, 16]. For example, Urban-SnowflakeNet [16] utilizes a deep learning network to extract features from preprocessed rooftop point clouds. It reconstructs building point clouds to aid in DSM rectification and TDOM generation, thereby effectively reducing distortions at building edges in urban scenes. However, these methods still rely on a combination of vision and LiDAR, requiring high-quality LiDAR point clouds and potentially failing in certain scenarios.
@@ 2.2 Deep Learning-based Methods
2.3 Differentiable Rendering-based Methods
The new view synthesis method based on differentiable volume rendering has significantly improved the quality of reconstructed images, providing a novel approach for pure vision-based TDOM generation. By integrating differentiable volume rendering, NeRF[7] trains an implicit neural radiance field, achieving remarkable-fidelity and highly continuous scene reconstruction. Many speed-optimized variants of NeRF have been applied to TDOM rendering, achieving image quality on par with traditional approaches [9].
@@ 2.3 Differentiable Rendering-based Methods
3 Method
The entire TDOM production process is shown in Figure 1. First, we initiate the point cloud as 2D Gaussians with original 2DGS training strategy presented in Subsubsection 3.1. The orthographic camera model and the depth map generation method are described respectively in Subsubsection 3.2 and Subsubsection 3.3. For the sake of training large-scale scenes, we employ a divide-and-conquer approach, and the seamlessly merged 2D Gaussians are rendered using Batch Rasterization to obtain high-resolution TDOM ,which is detailed in Subsubsection 3.4.
3.1 Preliminaries
@@
2DGS proposes flattening 3D Gaussian ellipsoids into 2D Gaussian elliptical disks. The planar-like shape of 2D Gaussian not only aligns more closely with the needs of surface reconstruction but also resolves the inconsistency in depth perception in 3DGS. As illustrated in Figure 2, this inconsistency arises when we observe Gaussian ellipsoids from different views, and the actual observed plane also varies. 2DGS constructs a local coordinate system for each 2D Gaussian. Any point $\mathbf{P}(u,v)$ in its local coordinate system can be transformed into its corresponding world coordinates with Gaussian center $\mu$, two principal tangential vectors $(\mathbf{t}_{u},\mathbf{t}_{v})$ and the scaling factors $(s_{u},s_{v})$:
@@ 3.2 Orthographic Projection of 2DGS
To obtain an orthographic image, we modify the projection method for the centers of the Gaussian primitives to orthographic projection. Shown as Figure 4, in the orthographic camera model, all rays are parallel to the optical axis of the camera. These two camera models are both demonstrated in Figure 3. Orthographic projection defines a square frustum as the visible space, and the projection matrix is defined as follows:
@@
3.3 Orthographic Depth Maps
@@
Since 2D Gaussian primitives are elliptical disks, when observed from a specific viewpoint, the primitives can degenerate from surfaces into lines, which is a degeneracy issue present in 2DGS. Shown as Figure 4, When the Gaussian value $G_{i}(\mathbf{x})$ at the ray-splat intersection is too small, we create a standard 2D Gaussian distribution $g_{i}^{2D}(\mathbf{x})$, using the projection of the Gaussian primitive center on the image plane as the mean. By comparing the values at the intersections of the ray with the two Gaussians, when $g_{i}^{2D}(\mathbf{x})$ is greater than $G_{i}(\mathbf{x})$, the 2D Gaussian for the current view is considered degenerate, and $d_{i}$ is taken as the z-axis coordinate of the Gaussian center in view space.
3.4 High-Resolution TDOM Rendering Strategy
@@ 4 Experiment
4.1 Experiments setup
Dataset: We selected several scenes with varying scales and characteristics from the NPU DroneMap Dataset[19], where the largest scene covers an area of 1.958 km². The chosen scenes include diverse land use types, including urban high-rise buildings, rural cottages, farmland, highways, and water bodies. The dataset was created by Bu et al., recorded using a custom-built hexacopter equipped with a GoPro Hero3+ camera and a DJI Phantom3. Each image has a 1920×1080 resolution. Moreover, we have also chosen a sample dataset provided by Agisoft [20], recorded by Canon DIGITAL IXUS 120 IS, which encompasses ground control point (GCP) information, to conduct a quantitative assessment of the TDOM’s quality. The panoramic images generated by our method are shown in Figure 5.
Implementation: The initial point cloud and camera poses obtained from COLMAP sparse reconstruction first undergo Manhattan alignment to align the scene’s bounding box with the coordinate axes. Each scene is divided into 8 partitions following the progressive partition strategy. Training is conducted in two batches across 4 NVIDIA GeForce RTX 3090 GPUs for 60,000 iterations, during this period, the attributes of the 2D Gaussians are continuously optimized. After training, the gaussians are filtered according to their central positions using bounding boxes from camera-based scene partitioning. The average training time for a scene is around 7 hours. Soon, the processed points will be stitched and deduplicated for the subsequent rendering. We use the orthographic camera model mentioned in Subsection3.2 to render orthophotos. The complete image is divided into 2$\times$2 tiles and rendered one by one. When each pixel is rendered in the GPU unit, the orthographic depth is also calculated.
4.2 TDOM Evaluation
@@
  • Buildings: In terms of high-quality TDOM, the buildings’ edges must be continuous and sharp, and their facades must not be visible. Our method excellently reconstructs the building boundaries and their occlusion relationships. In Figure 6(a) and Figure 6(b), it can be observed that other methods generally exist visible building facades, with Map2DFusion showing the most severe issues due to direct photo mosaicking without image correction. Additionally, ContextCapture has blurred edges, Metashape displays noticeable irregular boundaries, and Pix4Dmapper suffers from misaligned building edges.
  • Thin structures: Figure 6(c) and Figure 6(d) respectively demonstrate the restoration effects of the power tower and the air conditioner guardrail. Metashape consistently blurs and granulates scene information; the supports of the power tower become thickened and distorted in Map2DFusion, and are diminished in ContextCapture. The railings of the guardrail are significantly misaligned in all four software programs. In our method, these thin structures are clearer, and the details are well restored even under conditions of strong light and shadow contrast.
Table 1: Absolute distance error comparison between Metashape, ContextCapture, Pix4Dmapper and our method.
Pairs Absolute Error(m)$\downarrow$
Metashape ContextCapture Pix4Dmapper Ours
101-202 0.876 0.601 0.542 0.716
102-203 0.942 1.257 1.071 0.717
202-205 1.708 1.414 1.019 1.057
203-206 0.968 1.339 0.306 0.326
102-206 1.142 1.103 0.373 0.267
202-206 1.251 1.456 2.786 2.736
In the quantitative evaluation, the absolute distance error between two GCPs is applied as a criterion to evaluate the accuracy of the TDOM. We calculate the distances between the GCPs using both their true coordinate values and the coordinates extracted from the local coordinate system of the TDOM, while the true horizontal distances between the GCPs are computed using the Haversine formula. By aligning one pair of GCPs, we obtain a scaling factor, and then calculate the error between the aligned GCP distances from the TDOM and their true values. We compare our results with those from three commercial software solutions: Metashape, ContextCapture, and Pix4Dmapper. All the methods aligned the pair of control points GCP-101 and GCP-102. In our method, the minimum absolute error for GCP pairs is 0.267m, and the maximum absolute error is 2.736m. Our method maintains errors within 1.0m in the majority of cases and demonstrates comparable or superior accuracy relative to commercial software across most point pairs. Specific data can be found in Table 1.
@@ 4.3 Depth Map Evaluation
In the qualitative experiments, we replace the Red channel of the orthophoto with the depth map to facilitate an intuitive visualization of the correspondence between the depth information and the image. Additionally, we use the Canny edge detection algorithm [24] to extract building edges from the depth map and overlay them onto our generated TDOM to observe the alignment of the edges. As illustrated in Figure 7(1), The deeper red hue in the pixel indicates the higher height of the relevant surface object relative to the ground. The red areas predominantly cover the buildings on the surface, which aligns with real-world logic. In Figure 7(2), the edges extracted from the depth map distinctly outline the boundaries of the structures within the TDOM, as well as some of the taller trees on the terrain. Our depth map adds a third dimension of height information to the TDOM, which is beneficial for subsequent tasks such as image detection and semantic segmentation in TDOM.
5 Conclusion